# Robustness of timelike circular orbit topology against particle spin

### v1

@@ IV.3 Asymptotically dS black holes
indicating at least one unstable TCO for fixed $L$ and $s$.
@@
Set $M=1$, $\mathcal{M}=0.5$ and $l=10$. The graph of the function $f(r)$ is shown in Fig. 14, with $r_{h}\approx 2.09149$ and $r_{c}\approx 8.78885$.
@@
For $L=0.3$ and $1\leq s\leq 1$, the condition $2r-f^{\prime}s^{2}|_{r_{h}}>0$ satisfies. One unstable TCO exist (Fig. 11).
Figure 15: The graph of the potential function (56) for $L=0.3$ and $1\leq s\leq 1$. The red points and the blue boxes represent the maximum values and the minimum values of (56) for a given value of $s$, respectively.
@@ V Conclusions and discussions
In this work, we have applied a topological approach to study the circular orbits of spinning test particles in static, spherically symmetric spacetimes. By constructing an auxiliary potential and a vector field $\phi^{a}$ on the equatorial plane, we computed the topological winding number $W$ to classify the existence and stability of TCOs in various black hole spacetimes. Our main conclusions are:
@@
    Between Two Neighboring Horizons: In regions between two horizons (e.g., inner and outer black hole horizons, or between black hole and cosmological horizons in dS spacetime), we find W=1W=-1, indicating at least one unstable TCO for any fixed $L$ and $s$. This generalizes the result for non-spinning particles to spinning ones.
@@
    Outside the Outermost Horizon: In asymptotically flat and AdS spacetimes, we find $W=0$, meaning TCOs must occur in stable-unstable pairs or be absent. Numerical examples in Schwarzschild and Sch-AdS spacetimes confirm this. While the behavior of effective potentials for spinning particles and the existence of circular orbits have been studied since the pioneering work of Tod, de Felice, and Calvani Tod:1976ud , our topological approach reveals invariant properties ($W=0$) that are robust against the inclusion of spin.
  • (3).
    Independence of Spin Orientation: The topological number $W$ is independent of the spin magnitude $s$ and orientation (co- or counter-rotating). Spin affects the location and energy of orbits but not the global topological structure.
This work demonstrates the power of topological methods in classifying circular orbits, even for spinning particles. The invariance of $W$ under spin perturbations suggests a deep robustness in the orbital structure dictated by the background metric. This finding has important implications for gravitational wave astronomy, particularly for modeling extreme mass-ratio inspirals where the secondary object may possess significant spin.
Future directions include extending this analysis to stationary, axisymmetric spacetimes (e.g., Kerr black holes), where the interplay between black hole spin and particle spin may yield richer topological structures. Additionally, studying different spin-supplementary conditions could provide further insights into the gauge dependence of these results. From an astrophysical perspective, it would be valuable to explore the implications of these topological invariants for the gravitational-wave emission from spinning binaries and the formation of accretion disks around black holes.

### v2

@@ IV.3 Asymptotically dS black holes
indicating at least one unstable TCO for fixed $L$ and $s$. The cosmological horizon in dS spacetime introduces a natural boundary that modifies the topological classification. The guaranteed unstable TCO between the black hole and cosmological horizons may have implications for the global structure of orbits in our expanding universe, particularly for supermassive black holes where this region is astrophysically accessible.
@@
Set $M=1$, $\mathcal{M}=0.5$, $l=10$ and $1\leq s\leq 1$. The graph of the function $f(r)$ is shown in Fig. 16, with $r_{h}\approx 2.09149$ and $r_{c}\approx 8.78885$.
@@
For $L=0.3$, an unstable TCO exists for any value of $s$ (see Fig. 17).
Figure 17: The graph of the potential function (56) for $L=0.3$. The red points and the blue boxes represent the maximum values and the minimum values of (56) for a given value of $s$, respectively.
@@ V Conclusions and discussions
n this work, we demonstrate the power of topological methods in classifying circular orbits, even for spinning particles. The invariance of $W$ under spin perturbations suggests a deep robustness in the orbital structure dictated by the background metric. This finding has important implications for gravitational wave astronomy, particularly for modeling extreme mass-ratio inspirals where the secondary object may possess significant spin. Our main conclusions are:
@@
    In regions between two neighboring horizons—such as the inner and outer black-hole horizons or the black-hole and cosmological horizons of de-Sitter spacetime—we obtain W=1W=-1, guaranteeing at least one unstable TCO for every fixed pair $(L,s)$. Outside the outermost horizon, by contrast, asymptotically flat and AdS geometries yield $W=0$, so TCOs must either emerge in stable–unstable pairs or be absent altogether. Numerical tests in Schwarzschild-dS, Schwarzschild, and Schwarzschild-AdS backgrounds corroborate both statements, extending the earlier non-spinning-particle results to the spinning case.
@@
    The topological number $W$ is independent of the spin magnitude $s$ and orientation (co- or counter-rotating). Although the spin parameter $s$ indeed affects the dynamical behavior of TCOs, it does not change their global topological properties.
The spin-independence of the topological winding number implies that the fundamental orbital structure—such as the guaranteed existence of unstable TCOs between horizons or the pairing of stable and unstable TCOs outside—is a property of the spacetime itself, not of the particle’s internal composition. This universality underpins the robustness of gravitational wave signatures from spinning secondaries in EMRIs, suggesting that certain waveform features are more generic than previously assumed.
While our analysis demonstrates the robustness of topological invariants against spin perturbations, several limitations warrant discussion. The Tulczyjew spin-supplementary condition, though widely used, represents one particular choice for closing the MPD equations. Different supplementary conditions may yield quantitatively different orbital parameters, though we expect the topological invariants to remain unchanged due to their fundamental geometric nature.
Additionally, our restriction to static, spherically symmetric spacetimes excludes potentially interesting effects in rotating backgrounds. In Kerr spacetime, the interplay between black hole spin and particle spin may introduce new topological features that merit investigation. The small-spin approximation ($|s|<<M$) employed here ensures the validity of our analysis but leaves open the question of whether similar topological robustness persists for ultra-spinning particles.
Future work should also consider the astrophysical implications of these topological invariants for gravitational wave detection. The guaranteed existence of certain orbital types (e.g., unstable TCOs between horizons) may leave imprints in the gravitational wave signals from extreme mass-ratio inspirals, potentially providing new observational tests of strong-field gravity.